The lowest point in Death Valley is 86 meters below sea level. A helicopter flying at an altitude of 41 meters above this point climbed 28 meters and then dropped 37 meters. At what altitude relative to sea level was it then flying? Write an addition problem using integers and solve.
Answer by Theo(13342)



sea level is at 0 meters.
the floor of death valley is at -86 meters relative to sea level.
the helicopter starts off 41 meters above -86 meters.
this means the helicopter starts off at -86 + 41 = -45 meters relative to sea level.
the helicopter rises 28 meters to be at -45 + 28 = -17 meters relative to sea level.
the helicopter drops 37 meters to be at -17 - 37 = -54 meters relative to sea level.

-54 meters relative to sea level is 54 meters below sea level.